5.) 3.1 mol of instant tea is mixed with 236.6 g of H2O at a temperature of 90°C. At 90°C the vapor pressure of H2O alone is 525.26 mmHg. Calculate the vapor pressure of the tea solution.
1
Expert's answer
2020-05-21T13:52:35-0400

"Psolution = \u03a7solvent*P0solvent."

H2O moles=236.6/18 =13.14moles

Xsolvent =(13.14)/(13.14+3.1)= 0.809

P0solvent= 525.26mmHg

Psolution=0.809* 525.26= 424.94mmHg